A particle of mass M, initially at rest decays into two particles with rest masses m1 and m2 respectively. Show that the total energy of the mass m1 is

E1= c^(2) [M^(2)+ m1^(2) - m2^(2)] / 2M

c= speed of light
1
Expert's answer
2020-07-09T10:33:43-0400

Let's introduce 4 - vector "p_\mu = (\frac{E}{c},\vec p)", "p^\mu p_\mu = \left(\frac{E}{c}\right)^2-(\vec p)^2 = m^2c^2"

"p^{1}_\mu = (\frac{E_1}{c},\vec p_1) 
;p^{2}_\mu =(\frac{E_2}{c},\vec p_2);P_\mu =(\frac{E}{c},\vec P)"

where "P_\mu" - 4 - vector of momentum of the body "M"

"p^1_\mu, p^2_\mu" - 4 - vector of momentum of the "m_1, m_2"

We know that "\vec P = 0". Also we know that momentum and energy conserve laws are done in this case. So, we have the system of equations:

"\vec P = \vec p^1 + \vec p^2 = 0;"

"\left(\frac{E}{c}\right)^2 = M^2c^2"

"\left(\frac{E_1}{c}\right)^2-(\vec p^1)^2 = m_1^2c^2"

"\left(\frac{E_2}{c}\right)^2-(\vec p^2)^2 = m_2^2c^2"

We can solve this system relatively by the energies "E_1,E_2" and get the answer on this question.

"E_1 = c^2\frac{M^2+m^2_1-m^{2}_2}{2M}"